# iCost: A Novel Instance Complexity Based Cost-Sensitive Learning Framework for Imbalanced Classification

### v1

@@ Abstract
Class imbalance in data presents significant challenges for classification tasks. It is fairly common and requires careful handling to obtain desirable performance. Traditional classification algorithms become biased toward the majority class. One way to alleviate the scenario is to make the classifiers cost-sensitive. This is achieved by assigning a higher misclassification cost to minority-class instances. One issue with this implementation is that all the minority-class instances are treated equally, and assigned with the same penalty value. However, the learning difficulties of all the instances are not the same. Instances that are located near the decision boundary are harder to classify, whereas those further away are easier. Without taking into consideration the instance complexity and naively weighting all the minority-class samples uniformly, results in an unwarranted bias and consequently, a higher number of misclassifications of the majority-class instances. This is undesirable and to overcome the situation, we propose a novel instance complexity-based cost-sensitive approach in this study. We first categorize all the minority-class instances based on their difficulty level and then the instances are penalized accordingly. This ensures a more equitable instance weighting and prevents excessive penalization. The performance of the proposed approach is tested on 66 imbalanced datasets against the traditional cost-sensitive learning frameworks and a significant improvement in performance is noticeable, demonstrating the effectiveness of our method.
Index Terms:
@@
Cost-Sensitive Learning, Imbalanced Classification, Support Vector Machine, Data difficulty factors, Scikit-Learn
I Introduction
@@
When the class distribution in the dataset is uneven, with one class (the majority class) significantly outnumbering the other (the minority class), it is referred to as imbalanced data [1]. This type of data is frequently encountered in different applications such as medical diagnosis, fraud detection, spam detection, etc. [2]. The class imbalance can pose significant challenges for standard machine learning (ML) algorithms, which typically assume that the classes are balanced. Consequently, ML classifiers produce biased performance towards the majority class. Imbalanced learning is a critical area in ML that requires specialized techniques to ensure that models are effective and fair, especially in applications where the cost of misclassifying minority instances can be dire. As such, the imbalanced domain has caught a lot of attention from researchers and different approaches have been proposed to address the issue [3]. The techniques can be broadly classified into two categories: data-level approach and algorithmic-level approach.
@@
In the data-level approach, the original class distribution in the data is modified by adding new synthetic minority-class instances or eliminating samples from the majority class. The goal is to balance the class distribution in the data. Recent research suggests it is even more important to reduce the class overlapping in the process to obtain better performance [4]. On the other hand, in the algorithmic-level approach, the original classification algorithm is modified to adapt to the imbalanced domain scenario. This is achieved by changing the cost function to handle the class imbalance directly [5]. Higher misclassification costs are assigned to the minority class instances to make the algorithm more sensitive to those errors. During training, the model learns by trying to reduce the overall misclassification cost. Assigning higher weight to the minority-class misclassifications shifts the bias from the majority class. This way, the algorithm is made cost-sensitive (CS). This approach is classifier-dependent as different algorithms use different learning procedures. Both of these approaches perform almost equally well and are extensively used in different real-world applications [6, 7, 8].
@@
This study is focused on cost-sensitive learning. Here, a specific penalty is added to the misclassifications of the minority-class instances. Standard classifiers use a 0-1 loss function for calculating cost. This indicates a value of 0 for correct classifications and a value of 1 for incorrect classifications. This type of error-driven (ED) classifier assumes an even class distribution in the data. However, when the data is skewed, this approach does not fare well and fails to provide good sensitivity (accuracy of the minority class prediction). In many applications, correctly classifying the minority-class instances, which usually represent the positive cases, is more important. Therefore, the idea of the cost-driven classifier is introduced, where asymmetric misclassification cost is utilized. Assigning a higher misclassification cost to the minority-class instances compared to the majority-class forces the algorithm to put more priority on learning those instances correctly, reversing the bias. This approach works quite well when the data is imbalanced and has been incorporated into the implementation of different classification algorithms in scikit-learn and similar libraries such as xgboost [9].
[…]
TABLE I: Cost Matrix
@@
Figure 1: Effect of modifying weights on the decision function. Here, the size of points is proportional to its weight.
@@
In the case of the traditional CS approach, the assigned cost value is applied to all the minority class instances indiscriminately and this raises a major concern [10]. All minority class instances do not pose the same level of difficulty. Samples that are closer to the decision boundary have a higher chance of getting misclassified than those that are far away. The more difficult-to-learn samples should be penalized more heavily than the others. This instance-level difficulty characteristic has not been considered in previous literature and our study addresses this issue. For instance, a dataset has an IR of 100. Then all the minority-class instances will be penalized 100 times more strongly than any majority-class instances. This will create some unusual deformation of the decision boundary and consequently, cause a higher number of misclassifications of the majority class instances (lower specificity score) during testing. Assigning unnecessarily high penalties to the minority-class samples that are located far away from the decision boundary only biases the predictions toward the minority-class and fails to generalize well on the test data. To mitigate the situation, an instance-difficulty-based cost should be applied. We propose such an algorithm termed ”iCost: Instance complexity-based Cost-sensitive learning” in this study.
@@
Class overlapping is one of the major data difficulty factors that complicate the learning task when the data is imbalanced [4]. We try to address this issue in our proposed algorithm by first categorizing all the minority-class sample points based on their difficulty levels. We employ a neighborhood search algorithm to grade the samples into different levels. This is done based on the number of neighboring samples from the opposite class. A higher misclassification cost is assigned to the samples in the overlapping regions. This way, minority-class samples in the overlapped regions are prioritized significantly over the majority-class samples, leading to correct identifications of those hard-to-learn instances. On the other hand, samples that are in the safe zones surrounded by samples from the same class are marginally weighted, reducing their effect. The appropriate value for the penalty can be determined using grid search, evolutionary techniques, or similar approaches. This way, asymmetric cost is applied to different minority-class instances based on their complexity. This ensures a more appropriate distribution of weights among the minority-class instances. Thus, our proposed algorithm is able to provide improved prediction performance on a wide range of imbalanced datasets.
@@
The proposed technique is implemented by inheriting classifiers from the sklearn library, ensuring full compatibility with Sklearn’s CS classifiers. Different ways of categorization of the minority-class instances are possible and some of them have been included in our implementation. The performance of the proposed approach is evaluated on 66 different imbalanced datasets. Significant improvements in performance have been observed as compared to other CS approaches as well as popular sampling techniques.
@@
The rest of the article is organized as follows. Related works have been discussed in Section II. The proposed methodology is presented in Section III. In Section IV, we laid out the experimental setup. The performance measures are reported in Section V. The manuscript concludes with Section VI, with a summary of the work.
II Related Works
@@
A lot of different techniques have been proposed over the years to deal with imbalanced data [2]. However, very few of them address data-intrinsic characteristics [11, 12]. Several data difficulty factors have been identified which include class overlapping, small disjunct, noisy samples, etc [13]. It has been suggested that the overlapping between the classes, not the imbalance, is primarily responsible for complicating the learning task [14]. Recent literature emphasizes the importance of considering data difficulty factors to develop more robust methods for learning from imbalanced datasets [13]. Stefanowski proposed several approaches to tackle these data difficulty factors [15]. However, the study only considers data resampling techniques. In this study, we are primarily focused on CS approaches.
@@
Different classifiers such as SVM, ANN, and DTs are adapted to the CS framework [16, 17]. Their implementation is available in the popular sklearn library. In different medical datasets, class imbalance is prevalent and CS approaches are found to be quite successful in handling such scenarios [18]. Mohosheu et al. performed a detailed efficacy analysis of the performance of CS approaches as compared to data resampling techniques in their study [19]. The authors reported that the traditional CS approaches outperform undersampling and ensemble methods but cannot surpass popular oversampling techniques such as SMOTE or ADASYN. Hybridization between sampling and CS approaches is also possible and has been proposed by Newaz et al. [20]. The authors first resampled the data by generating synthetic samples using SMOTE and then used a weighted XGBoost classifier for training. The authors suggested that combining the techniques together allows for lower misclassification costs to be applied while also reducing the number of synthetic samples to be generated. Consequently, overfitting is reduced and better prediction performance can be obtained.
@@
Gan et al. proposed a sample distribution probability-based CS framework in their article [21]. Roychoudhuri et al. adapted the CS algorithm for time-series classification [22]. Zhou et al. extended the CS framework for multiclass imbalanced scenarios [23]. Other variations of CS approaches include MetaCost [24], a meta-learning algorithm that converts any given classifier into a CS classifier. The idea of example-dependent cost has also been proposed in previous literature [25, 26]. For instance, in credit scoring, a borrower’s credit risk is determined based on different factors which include their credit history and financial behaviors. These factors should be taken into consideration while weighting instances for predictive modeling [27]. However, such approaches are application-specific and do not generalize well for other datasets. A detailed review of different CS methods has been presented in this recent article [28]. None of these CS approaches considers instance-difficulty-based characteristics. This has been pointed out in some recent literature [4, 10, 13] and we address this issue in our study.
III Methodology
In this section, we discuss our proposed algorithm in detail. At first, we identify the K-nearest neighbors of each minority class sample. A k-value of 5 was utilized in this study. The nearest neighbors are computed using the Euclidean distance. Next, each minority class instance is categorized as follows:

### v2

@@ Abstract
Class imbalance in data presents significant challenges for classification tasks. It is fairly common and requires careful handling to obtain desirable performance. Traditional classification algorithms become biased toward the majority class. One way to alleviate the scenario is to make the classifiers cost-sensitive. This is achieved by assigning a higher misclassification cost to minority-class instances. One issue with this implementation is that all the minority-class instances are treated equally, and assigned with the same penalty value. However, the learning difficulties of all the instances are not the same. Instances that are located in the overlapping region or near the decision boundary are harder to classify, whereas those further away are easier. Without taking into consideration the instance complexity and naively weighting all the minority-class samples uniformly, results in an unwarranted bias and consequently, a higher number of misclassifications of the majority-class instances. This is undesirable and to overcome the situation, we propose a novel instance complexity-based cost-sensitive approach (termed ’iCost’) in this study. We first categorize all the minority-class instances based on their difficulty level and then the instances are penalized accordingly. This ensures a more equitable instance weighting and prevents excessive penalization. The performance of the proposed approach is tested on 65 binary and 10 multiclass imbalanced datasets against the traditional cost-sensitive learning frameworks. A significant improvement in performance has been observed, demonstrating the effectiveness of the proposed strategy.
keywords:
@@
Cost-Sensitive Learning , Imbalanced Classification , Multiclass classification , Data difficulty , Scikit-Learn , Weighted classifier
{highlights}
Cost-sensitive learning is a popular technique used in the imbalanced domain.
Higher misclassification costs are naively applied to all minority-class instances.
In the proposed approach, instances are penalized according to their complexity.
This ensures a plausible weighting mechanism, resulting in fewer misclassifications.
The iCost algorithm, written in Python, is fully compatible with the scikit-learn library.
1 Introduction
@@
When the class distribution in the dataset is uneven, with one class significantly outnumbering the other, it is referred to as imbalanced data [1]. Class imbalance in data is frequently encountered in different applications such as medical diagnosis, fraud detection, spam detection, etc. [2]. This can pose significant challenges for standard machine learning (ML) algorithms, which typically assume that the classes are balanced. Consequently, ML classifiers produce biased performance towards the majority class. Imbalanced learning is a critical area in ML that requires specialized techniques to ensure that models are effective and fair, especially in applications where the cost of misclassifying minority instances can be dire [3]. As such, the imbalanced domain has caught a lot of attention from researchers and different approaches have been proposed to address the issue [4]. The techniques can be broadly classified into two categories: data-level approach and algorithmic-level approach.
@@
In the data-level approach, the original class distribution in the data is modified by adding new synthetic minority-class instances or eliminating samples from the majority class [5]. The goal is to balance the class distribution in the data. Recent research suggests it is even more important to reduce the class overlapping in the process to obtain better performance [6]. On the other hand, in the algorithmic-level approach, the original classification algorithm is modified to adapt to the imbalanced domain scenario. This is achieved by changing the cost function to handle the class imbalance directly [7]. Higher misclassification costs are assigned to the minority class instances to make the algorithm more sensitive to those errors. During training, the model learns by trying to reduce the overall misclassification cost. Assigning higher weight to the minority-class misclassifications shifts the bias from the majority class. This way, the algorithm is made cost-sensitive (CS). This approach is classifier-dependent as different algorithms use different learning procedures. Both of these approaches perform almost equally well and are extensively used in different real-world applications [8, 9, 10, 11].
@@
This study is focused on cost-sensitive learning (CSL). Here, a specific penalty is added to the misclassifications of the minority-class instances. Standard classifiers use a 0-1 loss function for calculating cost. This indicates a value of 0 for correct classifications and a value of 1 for incorrect classifications. This type of error-driven (ED) classifier assumes an even class distribution in the data. However, when the data is skewed, this approach does not fare well and fails to achieve high sensitivity (accuracy of the minority class prediction). This happens because misclassifications of both majority and minority class instances are given equal weight. In many applications, correctly classifying the minority-class instances, which usually represent the positive cases, is more important [9]. Therefore, the idea of the cost-driven classifier is introduced, where asymmetric misclassification cost is utilized. Assigning a higher misclassification cost to the minority-class instances compared to the majority-class forces the algorithm to put more priority on learning those instances correctly, reversing the bias. This approach works quite well when the data is imbalanced and has been incorporated into the implementation of different classification algorithms in scikit-learn and similar libraries such as xgboost [12, 13].
Table 1: Cost Matrix
@@
[…]
The effect of modifying the weights of the instances on the decision boundary for the Support Vector Machine (SVM) classifier is illustrated in Fig 1. Assigning a higher weight to some instances forces the classifier to put more emphasis on getting those points right, deforming the original decision boundary. This often leads to misclassifications of some of the nearby unweighted samples, as can be observed from the figure. Therefore, instances should be weighted very carefully to avoid unusual predicaments.
@@
Figure 1: Effect of modifying weights on the decision boundary. Here, the size of points is proportional to its weight.
@@
In the case of the traditional CS approach, the assigned cost value is applied to all the minority class instances indiscriminately and this raises a major concern [14]. All minority class instances do not pose the same level of difficulty. Samples that are closer to the decision boundary have a higher chance of getting misclassified than those that are far away. The more difficult-to-learn samples should be penalized more heavily than the others. This instance-level difficulty characteristic has not been considered in previous literature and our study addresses this issue. For instance, let’s assume a dataset has an IR of 100. Then all the minority-class instances will be penalized 100 times more strongly than any majority-class instances. This will create some unusual deformation of the decision boundary and consequently, cause a higher number of misclassifications of the majority class instances (lower specificity score) during testing. Assigning unnecessarily high penalties to the minority-class samples that are located far away from the decision boundary only biases the predictions toward the minority-class. Consequently, some nearby majority-class instances get wrongly classified. Overall, the algorithm struggles to perform well on the test data. To mitigate the situation, an instance-difficulty-based cost should be applied. We propose such an algorithm termed ”iCost: Instance complexity-based Cost-sensitive learning” in this study.
Now, how to faithfully define the complexity of the instances remains an issue. Class overlapping between different class instances is usually identified as the most crucial factor that affects the classifier’s performance [15]. The presence of noisy samples in the data is another major concern. Small disjunct in the dataset also makes the learning task difficult for the algorithms [16]. These issues have been illustrated in Fig. 2 and Fig. 3. As we can see from figure 3, there are two minority-class instances completely surrounded by majority-class samples. These two samples are likely outliers as they are far away from other minority-class instances. However, in the traditional CS approach, these two noisy samples will also be penalized. This will obviously cause misclassifications of some of the nearby majority-class instances. Moreover, the model will fail to generalize well in unseen test data. As for figure 2, the instances in the overlapping region are more difficult to classify as they are surrounded by some opposite-class instances. These particular samples should be penalized more stringently than instances in the non-overlapping region. We consider all these parameters while defining instance complexity for our proposed algorithm.
Figure 2: Class overlapping between opposite class instances
Figure 3: Presence of noisy instance in the data (red circle marks the noisy minority-class instances). It also illustrates small disjunct in the data.
Researchers have proposed various ways of identifying overlapping regions. A taxonomy of different class overlap measures was presented by Santos et al. [17]. The authors grouped the approaches into four different categories: feature overlap, instance overlap, structural overlap, and multiresolution overlap. Each category of approach undertakes a different method for quantifying class overlap. For instance, maximum Fisher’s discriminant ratio (F1) is one of the most widely used feature-based measures to compute class overlap [18]. Clustering methods can also be utilized to determine the degree of overlap [19]. In this study, we used two different methodologies to define instance complexity. One is based on a neighborhood search algorithm. The other one uses concepts from graph theory by analyzing information on the internal structure of classes. These approaches allow us to take into consideration the class overlapping problem as well as the presence of noisy instance scenarios. In the future, we plan to incorporate a few other methods such as hypersphere coverage [20], and local set [21] to define instance complexity.
@@
In the first approach, a neighborhood search algorithm is employed to grade the samples into different levels. This is done based on the number of neighboring samples from the opposite class. A higher misclassification cost is assigned to the samples that are surrounded by more instances of the opposite class. This method also allows us to identify noisy instances in the data. This way, minority-class samples in the overlapped regions are prioritized significantly over the majority-class samples, leading to correct identifications of those hard-to-learn instances. On the other hand, samples that are in the safe zones surrounded by samples from the same class are marginally weighted, reducing their effect. In the second approach, a Minimum Spanning Tree (MST) is constructed over the data to determine structural overlap. The minority-class instances that are connected to the opposite class by an edge in the MST in considered boundary samples and therefore highly penalized compared to others. This way, asymmetric cost is applied to different minority-class instances based on their complexity. This ensures a more appropriate distribution of weights among the minority-class instances. This not only helps in reducing the majority-class misclassifications but also improves the performance of the minority-class. Thus, our proposed algorithm is able to provide improved prediction performance on a wide range of imbalanced datasets.
@@
The appropriate value for the penalty can be determined using grid search, evolutionary techniques, or similar approaches. Although these techniques are likely to enhance predictive performance, excessive optimization may overfit the data. The algorithm code is written in Python. It is implemented by inheriting classifiers from the sklearn library, ensuring full compatibility with sklearn’s CS classifiers. Different ways of categorization of the minority-class instances are possible and some of them have been included in our implementation. The performance of the proposed approach is evaluated on 65 binary imbalanced datasets. The algorithm is further extended for multiclass scenarios and tested on 10 multiclass imbalanced datasets. Significant performance improvements have been observed as compared to traditional CS approaches (employed in sklearn implementation) as well as some popular data resampling techniques. The results obtained affirm the advantage of the proposed approach.
@@
The rest of the article is organized as follows. Related works have been discussed in Section II. The proposed methodology is presented in Section III. In Section IV, we laid out the experimental setup. The performance measures are reported in Section V. The manuscript concludes with Section VI, with a summary of the work. For reproducibility and the benefit of the community, all codes and results are made available in the following GitHub repository: https://github.com/newaz-aa/iCost.
2 Related Works
@@
A lot of different techniques have been proposed over the years to deal with imbalanced data [22, 23]. However, very few of them address data-intrinsic characteristics [24]. Several data difficulty factors have been identified that are primarily responsible for making imbalanced classification so difficult [16]. This includes class overlapping, small disjuncts, noisy samples, and rarity of samples among others. It has been suggested that the overlapping between the classes, not the imbalance, is primarily responsible for complicating the learning task [25]. Recent literature emphasizes the importance of considering data difficulty factors to develop more robust methods for learning from imbalanced datasets [16]. Stefanowski proposed several approaches to tackle these data difficulty factors [26]. However, the study only considers data resampling techniques. In this study, we are primarily focused on CS approaches.
@@
Different classifiers such as SVM, Artificial Neural Network (ANN), and Decision Tree (DT) are adapted to the CS framework [27, 28]. Their implementation is available in the popular sklearn library. In different medical datasets, class imbalance is prevalent and CS approaches are found to be quite successful in handling such scenarios [29]. Mohosheu et al. performed a detailed efficacy analysis of the performance of CS approaches as compared to data resampling techniques in their study [30]. The authors reported that the traditional CS approaches outperform undersampling and ensemble methods but cannot surpass popular oversampling techniques such as SMOTE or ADASYN [31]. Hybridization between sampling and CS approaches is also possible [32, 33]. The authors first resampled the data by generating synthetic samples using SMOTE and then used a weighted XGBoost classifier for training. The authors suggested that combining the techniques allows for lower misclassification costs to be applied while also reducing the number of synthetic samples to be generated. Consequently, overfitting is reduced and better prediction performance can be obtained.
@@
Natarajan et al. explored two different ways to improve binary classification performance using CSL under noisy labels, focusing on real-world applications. The authors also reported that weighted LR and SVM classifiers are somewhat noise-tolerant. Gan et al. proposed a sample distribution probability-based CS framework in their article [34]. Roychoudhuri et al. adapted the CS algorithm for time-series classification [35]. Zhou et al. extended the CS framework for multiclass imbalanced scenarios [36]. Other variations of CS approaches include MetaCost [37], a meta-learning algorithm that converts any given classifier into a CS classifier. The idea of example-dependent cost has also been proposed in previous literature [38, 39]. For instance, in credit scoring, a borrower’s credit risk is determined based on different factors which include their credit history and financial behaviors. These factors should be taken into consideration while weighting instances for predictive modeling [40]. However, such approaches are application-specific and do not generalize well for other datasets. A detailed review of different CS methods has been presented in this recent article [41]. None of these CS approaches considers instance-difficulty-based characteristics. This has been pointed out in some recent literature [6, 14, 16] and we address this issue in our study.
3 Methodology
3.1 Cost Sensitive Classifier
In classification algorithms, a cost function is utilized for the learning process. The goal is to minimize the total cost associated with the predictions. Optimization algorithms such as stochastic gradient descent are utilized to reach the optimal point. This way, the model parameters are determined and the classifier learns to predict as accurately as possible. To give an example, for the LR classifier, log loss is calculated in the following way -
Log $\text{Log Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(h_{\theta}(x_{i}))+(1-y_{i})\log(1-h_{\theta}(x_{i}))\right]$
Here,
  • 1.
    N is the total number of samples
  • 2.
    $y_{i}$ is the true label of the ith sample (0 or 1)
  • 3.
    $h_{\theta}(x_{i})$ is the predicted probability that the ith sample belongs to class 1, as calculated by the classifier.
  • 4.
    $\theta$ is the model parameters.
The first term in the sum represents the cost for the positive class and the second term represents the cost for the negative class. Now, for the correct predictions, the cost becomes 0. As for the incorrect predictions, the cost is calculated for all misclassifications. Here, no distinction is made between the positive and negative instances. All the misclassifications are weighted equally. To make the classifier cost-sensitive, the log-loss formula can be modified by introducing different weights.
CS Log $\text{CS Log Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[C_{1}\cdot y_{i}\log(h_{\theta}(x_{i}))+C_{0}\cdot(1-y_{i})\log(1-h_{\theta}(x_{i}))\right]$
Here,
  • 1.
    $C_{1}$ is the cost of misclassifying a positive example.
  • 2.
    $C_{0}$ is the cost of misclassifying a negative example.
This formula allows adjusting the penalty for each class by tuning $C_{1}$ and $C_{0}$. This is the traditional CS approach where different weights are assigned to majority and minority class misclassifications. Usually, the $C_{0}$ value is kept fixed at 1 while the $C_{1}$ is assigned a different weight (typically, the IR of the dataset). This allows the algorithm to put more emphasis on correctly predicting the minority-class instances. Thus, modifying the learning process. While this approach certainly improves the model’s ability to correctly identify the minority-class instances, this usually comes at the cost of many misclassifications of the majority-class instances as they are assigned a lower weight. The trade-off becomes a defining factor here that how many misclassifications of the majority-class instances can be justified in exchange for a few correct predictions of minority-class instances. Although in certain applications such as medical diagnosis or fault detection, correctly identifying the positive cases is crucial, however, too many false positives reduce the reliability of the prediction framework.
In our proposed methodology, we try to address this critical but often neglected issue through an intuitive hypothesis. In the traditional CS framework, an underlying issue is ignored. All the instances in the dataset do not have the same difficulty of learning. Depending on their location in the feature space, some are easier to classify while others can be difficult. Instances that are closely located i.e., share similar feature attributes can be difficult to distinguish from one another. Depending on such characteristics, we propose to divide the instances into multiple categories. Then each category of samples is weighted differently in accordance with their complexity level. Mathematically, it can be represented as follows -
Modified CS Log $\text{Modified CS Log Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[C_{j(i)}\cdot y_{i}\log(h_{\theta}(x_{i}))+(1-y_{i})\log(1-h_{\theta}(x_{i}))\right]$
Here, $C_{j}(i)$ is the cost for the category j that instance i belongs to. The concept of categorization of the instances is discussed in the following section. One thing to consider here, only the minority-class instances (positive cases) are categorized and penalized with different weights. The majority-class instances are assigned equal weights, in line with the traditional cost-sensitive approach. While majority-class instances can also be categorized in a similar fashion, that would increase the complexity of the algorithm and is not explored in this article.
This concept is equally applicable to other classification algorithms such as SVM or DTs. For the SVM classifier, the cost (objective) function aims to find a hyperplane that maximizes the margin between classes while penalizing misclassified points. For a linear SVM, the objective function can be expressed as follows:
$\text{Minimize}\quad\frac{1}{2}||\mathbf{w}||^{2}+C\sum_{i=1}^{N}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))$
Here,
  • 1.
    w is the weight vector defining the hyperplane.
  • 2.
    $||w||^{2}$ is the regularization term that controls the margin width.
  • 3.
    C is the penalty parameter that controls the trade-off between maximizing the margin and minimizing classification errors.
  • 4.
    $y_{i}$ is the true label for instance i.
  • 5.
    $x_{i}$ is the feature vector for instance i.
  • 6.
    b is the bias term.
  • 7.
    $max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b)$ is the hinge loss, penalizing instances on the wrong side of the margin.
The function can be modified to better handle imbalanced data by penalizing errors on minority-class instances more heavily. The objective function for the CS-SVM classifier can be expressed as follows:
$\text{Minimize}\quad\frac{1}{2}||\mathbf{w}||^{2}+C_{+}\sum_{i\in\mathcal{P}}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))+C_{-}\sum_{i\in\mathcal{N}}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))$
Here,
  • 1.
    $C_{+}$ is the penalty parameter for the positive class (often the minority class).
  • 2.
    $C_{-}$ is the penalty parameter for the negative class (often the majority class).
  • 3.
    $\mathcal{P}$ and $\mathcal{N}$ represent the sets of positive and negative instances, respectively.
We can incorporate the proposed instance-based weighting formula into the equation in the following way -
$\text{Minimize}\quad\frac{1}{2}||\mathbf{w}||^{2}+C\sum_{i\in\mathcal{M}}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))+\sum_{j=1}^{k}C_{j}\sum_{i\in\mathcal{C}_{j}}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))$
Here,
  • 1.
    C is the penalty parameter for all majority-class instances, represented by set $\mathcal{M}$.
  • 2.
    $C_{j}$ is the penalty parameter specific to category j within the minority class.
  • 3.
    k is the number of categories in the minority class.
This way, the cost-sensitive learning framework of different classifiers can be adjusted to incorporate the instance-complexity-based weighting mechanism into the algorithm.
3.2 Instance Complexity
In this manuscript, we present two distinct approaches to defining instance complexity: one based on a neighborhood search algorithm and the other on a graph algorithm.
3.2.1 Neighborhood Search
When a minority-class instance is surrounded by majority-class samples, it naturally becomes difficult to classify them correctly. In the nearest neighbor rule, if there are more instances from the opposite class surrounding the given example, then it would be misclassified. We utilize this concept to identify hard-to-learn examples in the data. At first, we identify the K-nearest neighbors of each minority class sample. A k-value of 5 was utilized in this study. The nearest neighbors are computed using the Euclidean distance. Next, each minority-class instance is categorized as follows:
  • 1.